OpenStudy (anonymous):

Factor Completely: x^2 − 49

OpenStudy (anonymous):

use this rule \[a^2-b^2=(a-b)(a+b)\]

mathslover (mathslover):

It is just the reverse of the previous quest. \((a+b)(a-b) = a^2- b^2\) \(x^2-49\) is of the form \(a^2-b^2\)

OpenStudy (mayankdevnani):